im having a very hard time figuring out how to do the substitution method the problem that im stuck on is x=5+y then the second part of the equation is 2x-3y=7 but im stuck because i dont understand how to do it

Answer by rfer(16322)

x=5+y
2x-3y=7
---------
2(5+y)-3y=7
10+2y-3y=7
-y=7-10
-y=-3
y=3
-------------
2x-3*3=7
2x-9=7
2x=7+9
2x=16
x=8
----------
proof
x=5+y
8=5+3